# RobustExplain: Evaluating Robustness of LLM-Based Explanation Agents for Recommendation

### v1

@@ 2.1. Explainable Recommendation and LLM-Based Explanation Agents
More recently, Large Language Models (LLMs) have emerged as powerful explanation generators due to their ability to produce fluent, context-aware, and personalized natural-language justifications (Wu et al., 2024; Petruzzelli et al., 2024). LLM-based explanation agents can reason over user interaction histories, articulate preference patterns, and describe item characteristics in a human-like manner, substantially improving explanation quality and user comprehension (Liu et al., 2024). These capabilities have enabled conversational and agent-driven recommender systems, where explanations are generated dynamically as part of interactive user experiences (Gao et al., 2023; Peng et al., 2025).
@@ 2.2. Robustness in Explainable AI and Noisy Recommender Systems
Robustness has been widely studied across machine learning domains, particularly in natural language processing and explainable AI. In NLP, prior work has proposed robustness-aware evaluation metrics and perturbation-based analyses to assess the stability of text generation models under lexical, syntactic, or semantic input changes (Liu et al., 2023; Zhang et al., 2020; Sellam et al., 2020). Several studies have examined the sensitivity of LLMs to prompt variations or distributional shifts (Wang and others, 2024; Zheng et al., 2024; Chandrasekaran and others, 2025), revealing that minor input changes can lead to substantial output variation. Robustness of explanations has also been explored through explanation-consistency finetuning and robustness-aware objectives, primarily in classification or prediction tasks with static feature inputs (Chen et al., 2025; Raina and others, 2025).
Separately, recommender system research has extensively investigated the impact of noisy user behavior on model performance. Real-world interaction data is affected by accidental clicks, shared accounts, delayed logging, missing metadata, and evolving preferences (Fan and others, 2022; Toledo et al., 2021). Prior work has studied robustness of sequential recommenders under training data perturbations and preference drift (Betello et al., 2024; Chen et al., 2024; Han et al., 2024; Zhu et al., 2024), as well as methods for denoising or exploiting noisy signals for preference learning (Li et al., 2024; Pang et al., 2024). These studies primarily focus on recommendation accuracy or ranking robustness, rather than the stability of natural-language explanations.

### v2

@@ 2.1. Explainable Recommendation and LLM-Based Explanation Agents
More recently, Large Language Models (LLMs) have emerged as powerful explanation generators due to their ability to produce fluent, context-aware, and personalized natural-language justifications (Wu et al., 2024; Petruzzelli et al., 2024). LLM-based explanation agents can reason over user interaction histories, articulate preference patterns, and describe item characteristics in a human-like manner, substantially improving explanation quality and user comprehension (Liu et al., 2024). Representative systems include CARTS (Chen et al., 2025a), which coordinates multiple agents for structured textual summarization, ARAG (Maragheh et al., 2025), which combines agentic retrieval with personalized recommendation, and VL-CLIP (Giahi et al., 2025), which enhances multimodal embeddings through LLM augmentation. These capabilities have enabled conversational and agent-driven recommender systems, where explanations are generated dynamically as part of interactive user experiences (Gao et al., 2023; Peng et al., 2025).
@@ 2.2. Robustness in Explainable AI and Noisy Recommender Systems
Robustness has been widely studied across machine learning domains, particularly in natural language processing and explainable AI. In NLP, prior work has proposed robustness-aware evaluation metrics and perturbation-based analyses to assess the stability of text generation models under lexical, syntactic, or semantic input changes (Liu et al., 2023; Zhang et al., 2020; Sellam et al., 2020). Several studies have examined the sensitivity of LLMs to prompt variations or distributional shifts (Wang and others, 2024; Zheng et al., 2024; Chandrasekaran and others, 2025), revealing that minor input changes can lead to substantial output variation. Robustness of explanations has also been explored through explanation-consistency finetuning and robustness-aware objectives, primarily in classification or prediction tasks with static feature inputs (Chen et al., 2025b; Raina and others, 2025).
Separately, recommender system research has extensively investigated the impact of noisy user behavior on model performance. Real-world interaction data is affected by accidental clicks, shared accounts, delayed logging, missing metadata, and evolving preferences (Fan and others, 2022; Toledo et al., 2021). Recent advances in generative recommendation such as GRACE (Ma et al., 2025) and location-aware prediction with LLMs (Sun et al., 2025) demonstrate the growing role of language models in recommendation, and multi-agent collaborative reasoning systems require adaptive resource allocation for efficient deployment (Zhang et al., 2025), yet robustness evaluation remains underexplored. Prior work has studied robustness of sequential recommenders under training data perturbations and preference drift (Betello et al., 2024; Chen et al., 2024; Han et al., 2024; Zhu et al., 2024), as well as methods for denoising or exploiting noisy signals for preference learning (Li et al., 2024; Pang et al., 2024). These studies primarily focus on recommendation accuracy or ranking robustness, rather than the stability of natural-language explanations.
@@ 5. Discussion
Cross-Metric Analysis. Our multi-dimensional evaluation reveals meaningful correlations between robustness metrics (Table 6). Semantic similarity and keyword stability show strong positive correlation (r=0.72), suggesting that explanations preserving overall meaning also tend to retain recommendation-relevant terms. Structural consistency correlates moderately with both semantic (r=0.58) and keyword (r=0.65) metrics, indicating that explanation reorganization often accompanies content changes. Length stability shows weak correlations with all other metrics (r=0.24–0.31), confirming that verbosity varies independently of content stability. These patterns suggest that robustness-aware optimization could focus primarily on semantic and keyword preservation, treating length normalization as a secondary concern.
Perturbation Type Sensitivity. Analysis across perturbation types (Table 3) reveals that drift perturbation achieves the highest robustness (0.541 at Level 1), while shuffle perturbation presents the greatest challenge. This pattern is consistent across all models: the 70B model achieves 0.652 semantic similarity under drift versus 0.632 under shuffle (Table 5). The difference is statistically significant ($p<0.05$), suggesting LLMs are more sensitive to temporal order disruption than gradual preference changes. Missing value perturbation shows intermediate robustness (0.529 average), indicating that LLMs can partially compensate for incomplete data by focusing on available interaction signals. Noise injection and behavior dilution fall between these extremes, with robustness scores of 0.528 and 0.522 respectively.
Metric-wise Robustness Patterns. Examining robustness across individual metrics reveals consistent patterns (Table 5). Length stability achieves the highest scores across all models and perturbation types (0.695–0.752), indicating that LLMs maintain relatively consistent explanation verbosity regardless of input perturbations. Semantic similarity follows (0.578–0.652), demonstrating moderate preservation of overall meaning. Keyword stability shows intermediate performance (0.455–0.524), reflecting partial retention of recommendation-relevant terms. Structural consistency scores lowest (0.355–0.415), suggesting that explanation organization is most susceptible to input variations. These patterns hold across all four models, indicating that metric-level robustness characteristics are model-agnostic. The large gap between length stability and structural consistency (approximately 0.35 absolute difference) suggests that future robustness improvements should prioritize structural preservation while leveraging the inherent stability of length characteristics.
Model Architecture Comparison. While our primary analysis focuses on model size effects, we also observe architecture-specific patterns between LLaMA and Qwen model families. At comparable parameter counts (7B–8B), LLaMA 3.1-8B and Qwen2.5-7B achieve nearly identical robustness scores (0.492), suggesting that robustness characteristics may be more dependent on model capacity than architecture design. However, within the Qwen family, the 14B model shows 5.5% improvement over the 7B variant, while the cross-family comparison between LLaMA 3.1-70B and Qwen2.5-14B reveals a 2.5% gap favoring the larger model. These observations suggest that scaling benefits for robustness may follow diminishing returns, with the largest improvements occurring in the 7B–14B range.